OpenStudy (anonymous):

Suppose that F(x)=∫1 to x f(t)dt, where f(t)=∫1 to t^2 sqrt(6+(u^4))/ (u) du Find F″(2).

OpenStudy (anonymous):

Woah...

OpenStudy (anonymous):

Let me write this out.

OpenStudy (anonymous):

Here's a pic of the problem if you want to see it

zepdrix (zepdrix):

Hmm this is a cool problem :3 thinking...

myininaya (myininaya):

Like wouldn't just be we have F'(x)=f(x) by fundamental theorem of calculus then differentiate that again F''(x)=f'(x) then find f'(t) using the other equation thingy

myininaya (myininaya):

you know also by fundamental theorem thingy

OpenStudy (anonymous):

I don't know, sorry. I'm taking Calc 1 this semester, so, I'm not great with integrals yet.

myininaya (myininaya):

do you understand what I mean @lawls

OpenStudy (anonymous):

i'm in calc 1 too lol

myininaya (myininaya):

you can find what from \[F(x)=\int\limits_{1}^{x}f(t) dt \text{ that } F'(x)=f(x) \text{ and from that } F''(x)=f'(x) \\ \text{ then you can also find } f'(x) \text{ from } f(x)=\int\limits_{1}^{x^2}\frac{\sqrt{6+u^4}}{u} du\]

zepdrix (zepdrix):

Ooo neato :O

OpenStudy (anonymous):

i'm not sure if this is right but is f'(x) = sqrt (6 + (x^2)^4) / x^2 * 2x ?

myininaya (myininaya):

that looks good to me

myininaya (myininaya):

and that is also F''(x) since F''(x)=f'(x)

OpenStudy (anonymous):

ok. now i just plug in 2?

myininaya (myininaya):

yeah

OpenStudy (anonymous):

i got 16.18641406

OpenStudy (anonymous):

thank you!!

myininaya (myininaya):

or exact form sqrt(262)